OpenStudy (rainbow_rocks03):

PLZ HELP ME I will medal+fan you for BEST response. 25% of a number is 80. What is 60% of the number? A. 192 B. 12 C. 53 D. 320

OpenStudy (rainbow_rocks03):

BRB

OpenStudy (kendricklamar2014):

Let the variable X be the "number". \[x * (25/100) = 80\] \[x = 80 / (25/100)\]

OpenStudy (kendricklamar2014):

So:\[x = 320\] \[320 * (60/100) = 192\]

OpenStudy (kendricklamar2014):

So the answer is: A. 192

OpenStudy (anonymous):

\[(80/0.25)0.6 = ? \]

OpenStudy (anonymous):

192 is the answer

OpenStudy (rainbow_rocks03):

thx for the help @KendrickLamar2014

OpenStudy (kendricklamar2014):

No Problem